OpenStudy (anonymous):

evaluate the equation. [sqrt(100+h)]-10/h with limit h->0

OpenStudy (amistre64):

is the post correct? or is that a sqrt(10) instead?

OpenStudy (anonymous):

rationalize the numerator by multplying top and bottom by \[\sqrt{100-h}+10\]

OpenStudy (freckles):

(a-b)(a+b)=a^2-b^2

OpenStudy (freckles):

\[(\sqrt{100-h}+10)(\sqrt{100-h}-10)=(\sqrt{100-h})^2-10^2 \] just like the formula

OpenStudy (paxpolaris):

\[\lim_{h \to0} {\sqrt{100+h}-10\over h}\]

OpenStudy (anonymous):

\[\lim_{h \to0} {\sqrt{100-h}-10\over h}\times \frac{\sqrt{100-h}+10}{\sqrt{100-h}+10}\] \[=\frac{100-h-100}{h(\sqrt{100-h}+10)}\]\[=\frac{-h}{h(\sqrt{100-h}+10)}\] cancel

OpenStudy (anonymous):

then replace \(x\) by \(10\)

OpenStudy (zarkon):

is it 100+h or 100-h?

OpenStudy (freckles):

lol @satellite73 copied me

OpenStudy (anonymous):

100+h

OpenStudy (freckles):

wait i copied him

OpenStudy (freckles):

you still do it in a very similar way instead of -h on top you have+h and the same for the -h on bottom you have +h

OpenStudy (anonymous):

but then how do i find the limit? @freckles

OpenStudy (freckles):

do you @satellite73 last step?

OpenStudy (freckles):

cancel out the factor h on top with the factor h on bottom

OpenStudy (freckles):

then plug in 0 for remaining h's

OpenStudy (freckles):

\[\lim_{h \to0} {\sqrt{100+h}-10\over h}\times \frac{\sqrt{100+h}+10}{\sqrt{100+h}+10} =\lim_{h \rightarrow 0}\frac{(100+h)-100}{h(\sqrt{100+h}+10)} \\=\lim_{h \rightarrow 0} \frac{h}{h(\sqrt{100+h}+10)} \] do you see the h/h is 1 thing?

OpenStudy (freckles):

now just replace h with 0 and use order of operations

OpenStudy (anonymous):

i got 1/20 is that correct?

OpenStudy (freckles):

yep because 10+10=20 :) so we have 1/20 gj

OpenStudy (anonymous):

@freckles thank you!

OpenStudy (freckles):

np